# Hierarchical Budget Policy Optimization for Adaptive Reasoning

### v2

@@ Abstract
Large reasoning models achieve remarkable performance through extensive chain-of-thought generation, yet exhibit significant computational inefficiency by applying uniform reasoning strategies regardless of problem complexity. We present Hierarchical Budget Policy Optimization (HBPO), a reinforcement learning framework that enables models to learn problem-specific reasoning depths without sacrificing capability. HBPO addresses the fundamental challenge of exploration space collapse in efficiency-oriented training, where penalties on long output length systematically bias models away from necessary long reasoning paths. Through hierarchical budget exploration, our approach partitions rollout samples into multiple subgroups with distinct token budgets, aiming to enable efficient resource allocation while preventing degradation of capability. We introduce differentiated reward mechanisms that create budget-aware incentives aligned with the complexity of the problem, allowing models to discover natural correspondences between task requirements and computational effort. Extensive experiments demonstrate that HBPO reduces average token usage by up to 60.6% while improving accuracy by 3.14% across four reasoning benchmarks. Unlike existing methods that impose external constraints or rely on discrete mode selection, HBPO exhibits emergent adaptive behavior where models automatically adjust reasoning depth based on problem complexity. Our results suggest that reasoning efficiency and capability are not inherently conflicting, and can be simultaneously optimized through appropriately structured hierarchical training that preserves exploration diversity.
@@ 1 Introduction
Advances in large reasoning models have led to impressive performance on complex reasoning tasks through chain-of-thought methodologies (OpenAI, 2024; DeepSeek-AI, 2025b). However, these models exhibit fundamental inefficiency: they generate unnecessarily long reasoning chains even for simple problems, sometimes consuming thousands of tokens for basic arithmetic (Chen et al., 2025; 2024). This phenomenon reveals a fundamental misalignment, as current reasoning models lack the ability to adapt their computational effort to the actual complexity of problems.
@@
Figure 1: HBPO prevents exploration space collapse through hierarchical budget exploration. While efficient reasoning methods progressively abandon long reasoning paths during training, and length control methods like L1 achieve efficiency through mechanical constraints, HBPO partitions the exploration space into budget-constrained hierarchies (512, 1024, 2048, 2560 tokens). This structure maintains reasoning diversity throughout training, enabling emergent adaptive behavior where models match computational resources to problem complexity. The result is superior accuracy with efficient token usage varying from hundreds of tokens on GSM8K to thousands on AIME25.
@@
Recent empirical findings challenge the conventional belief that longer reasoning always leads to better outcomes. Research shows that models can maintain competitive accuracy even without intermediate steps (Ma et al., 2025), and in some cases, shorter reasoning paths perform comparably or even better on simpler tasks (Li et al., 2025). This is further supported by stark variations in optimal reasoning lengths across tasks. For instance, L1 (Aggarwal & Welleck, 2025) achieves peak performance with  1,100 tokens on GSM8K, but requires over 3,000 tokens on OlympiadBench. Such heterogeneity highlights a key insight: the computational requirements for effective reasoning are inherently problem-dependent, yet current models apply uniform reasoning strategies regardless of task complexity.
@@
To address these inefficiencies, an increasing number of studies aim to improve the inference efficiency of reasoning models. Current approaches fall into two primary categories. Length-controlled methods directly constrain generation through explicit mechanisms: prompts like “think for n tokens” and corresponding length-control rewards in L1 (Aggarwal & Welleck, 2025); progressively limits on the model’s reasoning space during training in ThinkPrune (Hou et al., 2025); and Scalable Chain of Thoughts (Xu et al., 2025) enforces budget constraints through forced termination. Reward-based approaches incorporate length penalties into training objectives: HAPO (Huang et al., 2025) leverages history-aware optimization to track minimal sufficient reasoning lengths, while “Think When You Need” (Jiang et al., 2025) employs pairwise comparison rewards to trade off between quality and brevity. While effective at reducing token usage, these methods share a key limitation: they prioritize efficiency at the cost of accuracy, lacking mechanisms for models to decide autonomously when longer or shorter reasoning is appropriate.
@@
We identify two key challenges that hinder existing methods from achieving genuine reasoning efficiency. First, length penalties introduce systematic training biases that impair reasoning capabilities. In standard reinforcement learning settings (DeepSeek-AI, 2025b), correct solutions receive equal rewards regardless of length, allowing for unbiased exploration. However, length penalties disrupt this balance by consistently favoring shorter outputs, leading policies to gradually abandon long-reasoning strategies (Hou et al., 2025; Huang et al., 2025; Lou et al., 2025). Second, static efficiency constraints fail to capture the continuous nature of reasoning complexity. Even adaptive methods rely on coarse mechanisms, such as binary think/no-think decisions (Zhang et al., 2025a; Fang et al., 2025) or fixed confidence thresholds (Qiao et al., 2025), which overlook the nuanced relationship between problem characteristics and computational requirements.
@@
We propose Hierarchical Budget Policy Optimization (HBPO), illustrated in the bottom-left part of Figure 1, a reinforcement learning framework that enables models to learn problem-specific reasoning strategies while retaining their ability to perform complex reasoning. The core idea is to partition the exploration space into multiple budget-constrained subgroups, allowing models to preserve reasoning diversity and uncover natural alignments between problem characteristics and required computational effort. Specifically, HBPO employs a hierarchical sampling strategy that partitions rollout samples into subgroups, each governed by a distinct token budgets. We implement this by inserting length prompts (e.g., “I will answer the question within n tokens”) after the reasoning tag, thereby constructing multiple exploration spaces with budgets ranging from 512 to 2560 tokens. Unlike uniform sampling, this structure encourages the model to explore both concise and extended reasoning paths throughout training, effectively mitigating the systematic degradation of reasoning capabilities caused by global length penalties.
To guide efficient reasoning within each budget hierarchy, we introduce a piecewise reward function that combines the strengths of classical and cosine-shaped reward forms. Within the assigned budget, rewards are monotonically non-decreasing to preserve exploratory flexibility. Beyond the budget, cosine decay and length deviation penalties are applied to encourage the model to return to its designated exploration space. This design establishes differentiated incentives across subgroups: shorter budgets favor concise solutions with higher rewards, while longer budgets retain standard rewards for extended reasoning, enabling adaptive resource allocation in line with problem complexity.
Our evaluation on four reasoning benchmarks shows that HBPO achieves a superior trade-off between efficiency and accuracy compared to existing approaches, as the results shown in the right part of Figure 1. It consistently outperforms both length-controlled methods with hard constraints and adaptive methods based on discrete mode selection. More importantly, HBPO demonstrates genuinely adaptive behavior: it automatically adjusts token usage, ranging from hundreds for simple problems to thousands for complex reasoning tasks, indicating that the model has learned to allocate computational resources based on problem complexity.
@@ 3 Method
We present Hierarchical Budget Policy Optimization, as shown in Figure 2, which extends the Group Relative Policy Optimization (GRPO) (DeepSeek-AI, 2025b) framework to enable adaptive reasoning through structured exploration. The core innovation lies in partitioning the exploration space into budget-constrained hierarchies and designing differentiated reward mechanisms that preserve reasoning diversity. We first introduce the hierarchical rollout strategy (Section 3.1), then detail the budget-aware reward design (Section 3.2)), and finally describe the training procedure (Section 3.3)).
@@ 3.2.1 Intra-Budget Reward Function
$R(n_{\text{gen}}\mid b)=\begin{cases}f_{1}(n_{\text{gen}},b),&\text{if correct, }n_{\text{gen}}>b,\text{ and }n_{\text{gen}}\leq L_{\max}\\ f_{2}(b),&\text{if correct, }n_{\text{gen}}\leq b,\text{ and }n_{\text{gen}}\leq L_{\max}\\ 0,&\text{otherwise}\end{cases}$ (1)
@@ 3.2.2 Inter-Budget Reward Differentiation
1:Initial policy $\pi_{\theta_{0}}$, budget levels =$\mathcal{B}=\{b_{1},...,b_{k}\}$, learning rate $\eta$
2:for iteration $t=1,2,...,T$ do
@@ 4.1 Experimental Setup
We evaluate HBPO on mathematical reasoning tasks using the DeepScaleR dataset (Luo et al., 2025) for training, which comprises 40K high-quality mathematical problems from AIME, AMC, Omni-Math (Gao et al., 2025), and STILL (Min et al., 2024). We employ two base models: DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025a) and DeepScaleR-Preview-1.5B (Luo et al., 2025).
@@
We evaluate on four mathematical reasoning benchmarks of increasing difficulty: GSM8K (Cobbe et al., 2021), Math500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), and AIME25. Following standard practice (DeepSeek-AI, 2025b), we use temperature $T=0.6$, $top\_p=0.95$, and maximum context length of 32,768 tokens. We report pass@1 accuracy and average token usage under two evaluation settings: (1) natural reasoning where models freely determine their computational effort, and (2) efficiency prompting using “I will answer the question with minimal tokens” after <think> to guide models toward efficient responses.
@@
We compare against several state-of-the-art efficient reasoning methods: (1) TLMRE(Arora & Zanette, 2025), which adds length penalties to the RL objective with hyperparameter $\alpha$ controlling the penalty strength; (2) AdaptThink(Zhang et al., 2025a) and AutoThink(Tu et al., 2025), which enable binary think/no-think mode selection; (3) L1-Max(Aggarwal & Welleck, 2025), which uses two-stage RL with explicit length targets. These methods represent different approaches to reasoning efficiency: global penalties (TLMRE), discrete mode selection (AdaptThink, AutoThink), and explicit control (L1-Max).
@@ 4.2 Main Results
Table 1: Performance under natural reasoning conditions. Models freely allocate computational resources based on learned strategies.
@@
Table 2: Performance under efficiency prompting reasoning. Models are explicitly prompted to minimize token usage.
@@ 5 Analysis
Figure 3: Training Dynamics: Mean Token Length (solid Line) ± Standard Deviation (dotted line).
@@ 5.1 Analysis of Hierarchical Structure
Training dynamics demonstrate how hierarchical structure maintains exploration space.
Figure 3 shows the generation dynamics between HBPO and a single-budget baseline. While single-budget training converges to a narrower range of responses, HBPO exhibits different dynamics. The average generation length stabilizes around 1,400 tokens, accompanied by broader yet controlled variance — a direct consequence of its hierarchical structure. This sustained variability is essential, as it captures the preservation of exploration diversity, which is crucial in preventing degradation of reasoning capability.
@@
As results shown in Table 4 and Figure 4, traditional GRPO with cosine reward achieves some efficiency (average 1,150 tokens) but suffers significant accuracy degradation, particularly on complex tasks where it achieves only 23.3% on AIME25. The model learns to generate universally short responses regardless of problem requirements, a form of mode collapse that sacrifices capability for efficiency.
@@
Figure 4: Comparison of training dynamics and validation performance between cosine reward and budget-aware reward methods. (Left) Training token length evolution shows that cosine reward exhibits a sharp decline in early training stage(0-100 steps), while budget-aware reward maintains relatively consistent token lengths throughout training. (Middle) Validation accuracy demonstrates that budget-aware reward achieves sustained performance improvements with higher final accuracy, whereas cosine reward shows greater volatility and inferior performance. (Right) Validation token count reveals that budget-aware reward enables the model to gradually discover the optimal reasoning length, while cosine reward suffers from excessive compression from the early training stages, leading to suboptimal token generation.
HBPO encourages effective exploration during training by employing budget-aware rewards rather than uniform compression, enabling the model to identify optimal reasoning lengths for efficient inference more accurately.
@@ 5.2 Reasoning Pattern Analysis
L1-Max improves efficiency through uniform length control, maintaining nearly constant thinking proportions (90-92%) and keyword frequencies (29-32) across three datasets. This rigidity reveals mechanical optimization rather than intelligent adaptation. AutoThink attempts adaptive reasoning but exhibits problematic patterns: excessive thinking on simple problems (86% on GSM8K) and insufficient adjustment for complex ones. Moreover, AutoThink exhibits an average of 1.7 and 1.5 reasoning-related keywords per problem in the solution segments on the MATH500 and Olympiad benchmarks, indicating that reasoning processes leak into what should be direct answers.

### v3

@@ Abstract
Large reasoning models achieve remarkable performance through extensive chain-of-thought generation, yet they suffer from a critical inefficiency: applying uniformly extensive reasoning regardless of problem complexity. We present Hierarchical Budget Policy Optimization (HBPO), a reinforcement learning framework that enables models to learn problem-specific reasoning depths without sacrificing capability. Unlike existing approaches that impose rigid constraints or rely on discrete mode selection, HBPO partitions the exploration space into budget-constrained hierarchies (512-2560 tokens), each with differentiated reward structures that preserve both efficiency incentives and reasoning capabilities. This design addresses a fundamental challenge in efficient reasoning training: traditional length penalties systematically bias models away from necessary long reasoning paths, causing exploration space collapse. Through hierarchical sampling and budget-aware rewards, HBPO maintains exploration diversity while teaching models to recognize when extended deliberation is warranted. Extensive experiments demonstrate that HBPO reduces average token usage by up to 60.6% while improving accuracy by 3.14% across four reasoning benchmarks. Most notably, HBPO exhibits emergent adaptive behavior where models automatically adjust reasoning depth based on problem complexity. Our results suggest that reasoning efficiency and capability are not inherently conflicting, and can be simultaneously optimized through appropriately structured hierarchical training that preserves exploration diversity.
@@ 1 Introduction
Figure 1: HBPO provides budget-aware reward through hierarchical budget exploration, which enables fine-grained adaptive reasoning. While length-constrained methods use global constraint or length penalty, and discrete mode methods dichotomize problem difficulty, HBPO partitions the exploration space into budget-constrained hierarchies (512, 1024, 2048, 2560 tokens). This structure maintains reasoning diversity throughout training, enabling emergent adaptive behavior where models match computational resources to problem complexity.
Advances in large reasoning models have led to impressive performance on complex reasoning tasks through chain-of-thought methodologies (OpenAI, 2024; DeepSeek-AI, 2025). However, these models exhibit fundamental inefficiency: they generate unnecessarily long reasoning chains even for simple problems, sometimes consuming thousands of tokens for basic arithmetic (Chen et al., 2025; 2024). This phenomenon reveals a fundamental misalignment, as current reasoning models lack the ability to adapt their computational effort to the actual complexity of problems.
@@
Recent empirical findings challenge the conventional belief that longer reasoning always leads to better outcomes. Research shows that models can maintain competitive accuracy even without intermediate steps (Ma et al., 2025), and in some cases, shorter reasoning paths perform comparably or even better on simpler tasks (Li et al., 2025). This is further supported by stark variations in optimal reasoning lengths across tasks. For instance, L1 (Aggarwal & Welleck, 2025) achieves peak performance with $\sim$1,100 tokens on GSM8K, but requires over 3,000 tokens on OlympiadBench. Such heterogeneity highlights a key insight: the computational requirements for effective reasoning are inherently problem-dependent, yet current models apply uniform reasoning strategies regardless of task complexity.
@@
To address these inefficiencies, an increasing number of studies aim to improve the inference efficiency of reasoning models. Current approaches fall into two primary categories. Length-constrained methods directly constrain generation through explicit mechanisms or incorporate length penalties into training objectives: prompts like “think for n tokens” and corresponding length-control rewards in L1 (Aggarwal & Welleck, 2025); progressively limits on the model’s reasoning space during training in ThinkPrune (Hou et al., 2025); enforces budget constraints through forced termination in Scalable Chain of Thoughts (Xu et al., 2025); and HAPO (Huang et al., 2025) leverages history-aware optimization to track minimal sufficient reasoning lengths. Discrete mode methods dichotomize problem difficulty and omit the reasoning process for simple instances, which enables the model to operate in a think/no-think manner. Thinkless (Fang et al., 2025) first performs format training for mode switching via fine-tuning. AdaptThink (Zhang et al., 2025a) employs importance sampling to enable the model to switch between reasoning patterns. While effective at reducing token usage, these methods share a key limitation: they prioritize efficiency or mode selection at the cost of accuracy performance, lacking fine-grained mechanisms for models to autonomously decide appropriately efficient reasoning length.
@@
We identify two key challenges that hinder existing methods from achieving genuine reasoning efficiency. First, length penalties introduce systematic training biases that impair reasoning capabilities. In standard reinforcement learning settings (DeepSeek-AI, 2025), correct solutions receive equal rewards regardless of length, allowing for unbiased exploration. However, length penalties disrupt this balance by consistently favoring shorter outputs, leading policies to gradually abandon long-reasoning strategies (Hou et al., 2025; Huang et al., 2025; Lou et al., 2025). Second, static efficiency constraints fail to capture the continuous nature of reasoning complexity. Even adaptive methods rely on coarse mechanisms, such as binary think/no-think decisions (Zhang et al., 2025a; Fang et al., 2025) or fixed confidence thresholds (Qiao et al., 2025), which overlook the nuanced relationship between problem characteristics and computational requirements.
@@
We propose Hierarchical Budget Policy Optimization (HBPO) illustrated in Figure 1, a reinforcement learning framework that enables models to learn problem-specific reasoning strategies while retaining their ability to perform complex reasoning. The core idea is to partition the exploration space into multiple budget-constrained subgroups, allowing models to preserve reasoning diversity and uncover natural alignments between problem characteristics and required computational effort. Specifically, HBPO employs a hierarchical sampling strategy that partitions rollout samples into subgroups, each governed by a distinct token budgets. We implement this by inserting length prompts (e.g., “I will answer the question within n tokens”) after the reasoning tag, thereby constructing multiple exploration spaces with budgets ranging from 512 to 2560 tokens. Unlike uniform sampling, this structure encourages the model to explore both concise and extended reasoning paths throughout training, effectively mitigating the systematic degradation of reasoning capabilities caused by global length penalties.
To enable efficient reasoning within each budget hierarchy, we design a piecewise reward function with distinct behaviors inside and outside budget boundaries. Within the assigned budget, rewards are monotonically non-decreasing to preserve exploratory flexibility. Beyond the budget, cosine decay and length deviation penalties are applied to encourage the model to return to its designated exploration space. This creates a natural gradient of incentives: shorter budgets favor concise solutions with higher rewards, while longer budgets retain standard rewards for extended reasoning. The result is a reward landscape that teaches models not just to reason efficiently within constraints, but to recognize which constraint level matches the problem at hand.
HBPO achieves a superior accuracy-efficiency trade-off compared to existing methods on four reasoning benchmarks. Crucially, it exhibits adaptive behavior by dynamically allocating computational resources based on problem complexity. For example, on GSM8K, it uses only 670 tokens. On AIME25, it uses 5,606 tokens, representing a more than eightfold increase in token usage. In both cases, it improves accuracy by 2.2% and 8.9% compared to the base model DeepSeek-R1-Distill-Qwen-1.5B, demonstrating effective resource allocation.
@@ 3 Method
We present Hierarchical Budget Policy Optimization, as shown in Figure 2, which extends the Group Relative Policy Optimization (GRPO) (DeepSeek-AI, 2025) framework to enable adaptive reasoning through structured exploration. The core innovation lies in partitioning the exploration space into budget-constrained hierarchies and designing differentiated reward mechanisms that preserve reasoning diversity. We first introduce the hierarchical rollout strategy (Section 3.1), then detail the budget-aware reward design (Section 3.2)), and finally describe the training procedure (Section 3.3)).
@@ 3.2.1 Intra-Budget Reward Function
[…]
@@ 3.2.2 Inter-Budget Reward Differentiation
1:Initial policy $\pi_{\theta_{0}}$, budget levels $\mathcal{B}=\{b_{1},...,b_{k}\}$, learning rate $\eta$
2:for iteration t=1,2,,Tt=1,2,...,T do
@@ 4.1 Experimental Setup
We evaluate HBPO on mathematical reasoning tasks using the DeepScaleR dataset (Luo et al., 2025) for training, which comprises 40K high-quality mathematical problems from AIME, AMC, Omni-Math (Gao et al., 2025), and STILL (Min et al., 2024). We employ two base models: DeepSeek-R1-Distill-Qwen-1.5B (DeepSeek-AI, 2025) and DeepScaleR-Preview-1.5B (Luo et al., 2025).
@@
We evaluate on four mathematical reasoning benchmarks of increasing difficulty: GSM8K (Cobbe et al., 2021), Math500 (Lightman et al., 2023), OlympiadBench (He et al., 2024), and AIME25. Following standard practice (DeepSeek-AI, 2025), we use temperature $T=0.6$, top_p=0.95top\_p=0.95, and maximum context length of 32,768 tokens. We report pass@1 accuracy and average token usage under two evaluation settings: (1) natural reasoning where models freely determine their computational effort, and (2) efficiency prompting using “I will answer the question with minimal tokens” after <think> to guide models toward efficient responses.
@@
We compare against several state-of-the-art efficient reasoning methods: (1) global penalties: HAPO  (Huang et al., 2025) and TLMRE (Arora & Zanette, 2025) add length penalties to the RL objective; (2) explicit control: L1-Exact,L1-Max (Aggarwal & Welleck, 2025), E1 (Xu et al., 2025) and ThinkPrune (Hou et al., 2025)use RL with explicit length targets. (3) discrete mode selection: AdaptThink (Zhang et al., 2025a), AutoThink (Tu et al., 2025) AdaR1 (luoAdaR1LongCoTHybridCoT2025) and Thinkless (Fang et al., 2025) enable binary think/no-think mode selection.
@@ 4.2 Main Results
Table 1: Performance under natural reasoning setting. Bold indicates the best and underline indicates the second-best for each metric. HBPO achieves the best performance in terms of the accuracy-efficiency trade-off and exhibits adaptive behavior.
@@
Table 2: Performance under efficiency prompting setting. HBPO demonstrates robust performance compared to baseline models and the explicit length-controlled method L1, while effectively adhering to efficient prompting instructions.
@@ 5.1 Analysis of Hierarchical Structure
As results shown in Table 4, traditional GRPO with cosine reward achieves some efficiency (average 1,150 tokens) but suffers significant accuracy degradation, particularly on complex tasks where it achieves only 23.3% on AIME25. The model learns to generate universally short responses regardless of problem requirements, a form of mode collapse that sacrifices capability for efficiency.
@@
Figure 3: Training dynamics. (Left) Entropy Comparison of different hierarchical structures. (Right) Comparison of training dynamics and validation performance between cosine and budget-aware reward methods.
Figure 3 presents the training dynamics of entropy, mean generating length, and validation on the Math500 dataset, highlighting the advantages of hierarchical structures and budget-aware reward mechanism. HBPO (4-budget) setting significantly increases entropy throughout training, outperforming both the dual-budget and single-budget baselines. This suggests that a more fine-grained budget hierarchy encourages more diverse and effective exploration, thereby preventing exploration collapse. When comparing cosine reward to HBPO(budget-aware reward), the cosine reward leads to a sharp drop in generation length during the early training stages (steps 0–100), which results in excessive compression and poor generalization on the Math500 validation set. In contrast, HBPO maintains a stable average generation length of approximately 1,400 tokens. This stability stems from its hierarchical structure, which encourages effective exploration through budget-aware rewards rather than uniform compression. As a result, the model gradually discovers the most efficient reasoning length on the Math500 validation set during training and consistently improves its validation accuracy.
@@ 5.2 Reasoning Pattern Analysis
L1-Max improves efficiency through uniform length control, maintaining nearly constant thinking proportions (90-92%) and keyword frequencies (29-32) across three datasets. This rigidity reveals mechanical optimization rather than intelligent adaptation. AutoThink attempts adaptive reasoning but exhibits problematic patterns: excessive thinking on simple problems (1171 tokens on GSM8K) and insufficient adjustment for complex ones. Moreover, AutoThink exhibits an average of 1.8 and 4.0 reasoning-related keywords per problem in the solution segments on the MATH500 and Olympiad benchmarks, indicating that reasoning processes leak into what should be direct answers.